# How weak are weak factors? Uniform inference for signal strength in signal plus noise models

### v1

@@ Abstract.
The paper analyzes four classical signal-plus-noise models: the factor model, spiked sample covariance matrices, the sum of a Wigner matrix and a low-rank perturbation, and canonical correlation analysis with low-rank dependencies. The objective is to construct confidence intervals for the signal strength that are uniformly valid across all regimes – strong, weak, and critical signals. We demonstrate that traditional Gaussian approximations fail in the critical regime. Instead, we introduce a universal transitional distribution that enables valid inference across the entire spectrum of signal strengths. The approach is illustrated through applications in macroeconomics and finance.
Gorin’s work was partially supported by NSF grant DMS-2246449. Sodin’s work was partially supported by a Philip Leverhulme Prize (PLP-2020-064) and by the Morris Belkin Visiting Professor Program at the Weizmann Institute of Science.
@@ 1.1. Motivation
In the modern era researchers increasingly have access to high-dimensional data sets across a wide range of fields. These data sets are inevitably contaminated by various forms of error and noise, making the separation of meaningful structure from background noise a central challenge. To address this analysts commonly employ dimension-reduction techniques. The two dominant approaches are low-rank methods, which assume that the underlying signal lies in a lower-dimensional subspace, and sparsity-based methods, which assume that only a small subset of variables or parameters are truly relevant, i.e., non zero. This paper adopts the low-rank perspective. For a discussion of settings where this assumption is appropriate, we refer to Udell and Townsend (2019), Giannone et al. (2021), and Thibeault et al. (2024). In particular, Giannone et al. (2021) argue that numerous data sets in macroeconomics, microeconomics, and finance exhibit dense, rather than sparse, structures.
@@
The feasibility of consistently estimating the signal component $LF^{\mathsf{T}}$ from the observed data $X$ hinges on the strength of the signal, which can be quantified by the singular values of $LF^{\mathsf{T}}$. When these singular values are large, the signal is strong and estimation is reliable, as can be directly predicted from the form of (1.1). As the signal weakens, the data $X$ becomes less informative, and below a certain critical threshold, accurate recovery becomes impossible. The relationships between the strength of the signal and feasibility of reconstruction of $LF^{\mathsf{T}}$ have been rigorously analyzed in a number of studies, see, e.g., Stock and Watson (2002); Bai and Ng (2002); Bai (2003); Paul (2007); Onatski (2012); Johnstone and Paul (2018); Bai and Ng (2023); Fan et al. (2024); Barigozzi and Hallin (2024) and references therein.
@@ 1.2. Models and results
We analyze four classical high-dimensional statistical models: the factor model (1.1), spiked sample covariance, the spiked Wigner model, and spiked canonical correlations. These models correspond to three fundamental ensembles from random matrix theory: the Laguerre/Wishart ensemble for the first two, the Hermite/Gaussian/Wigner ensemble for the third, and the Jacobi ensemble for the fourth. Each of these models can be viewed as an instance of the signal-plus-noise framework – also known as spiked random matrices, a term originating with Johnstone (2001) – in which a low-rank signal matrix is embedded in a high-dimensional noisy environment. The goal is to detect and quantify the embedded signal.
@@
Our analysis is based on spectral methods, whereby signal strength is inferred from the eigenvalues of certain model-specific matrices. In all four setups a well-documented phase transition phenomenon arises: the signal strength can be consistently estimated (in the high-dimensional asymptotic regime with proportional growth of data dimensions) only when it exceeds a critical threshold, see Jones et al. (1978), Baik and Silverstein (2006), Onatski (2012), Bao et al. (2019) and more references in Section 2. When the signal strength falls below the threshold, only partial probabilistic information, such as asymptotics of the likelihood ratio test can be recovered about the signal, but reliable point estimation becomes impossible, see, e.g. Onatski et al. (2013, 2014); Dobriban (2017); Johnstone and Onatski (2020); El Alaoui et al. (2020). The intermediate regime, where the signal strength is close to the threshold, is typically referred to as the “critical” regime. This regime is particularly challenging for inference.
@@
Our paper fills this gap by proposing a general procedure for constructing confidence intervals for signal strength. Remarkably, across all four models we study, the confidence intervals are characterized by a common limiting (stochastic) object, which we call the Airy–Green function and denote $\mathcal{G}(w)$. Our main contributions are: a rigorous construction of this function, a unified set of theorems linking it to the four canonical models, and tabulated confidence intervals based on $\mathcal{G}(w)$. The only model-specific components are a set of scaling constants, which we provide explicitly for each setting.
@@ 1.3. Econometrics and statistics contributions
In economics and finance it has long been observed that many data sets contain factors that are either non-informative or far from strong – see e.g. Giglio et al. (2023) and Kim et al. (2024) for overviews and extensive references. This concern is especially apparent in the vast “factor zoo” of potential variables proposed to explain stock returns. This empirical reality has motivated a line of theoretical research focused on inference for weaker factors. Broadly speaking, factors can be classified by their strength into three categories: strong (as in, e.g. Bai and Ng (2002); Stock and Watson (2002)), semi-strong (as in, e.g. Bai and Ng (2023); Fan et al. (2024)), and weak [^2] What we call semi-strong factors are sometimes referred to as weak, while weak factors may be termed weakly influential or extremely weak. (as in, e.g., Onatski (2012)). The literature also includes statistical procedures for testing and distinguishing between these types of factors (see, in particular, Kim et al. (2024)). Over the past decades a growing body of research has focused specifically on factor strength, including contributions by Chudik et al. (2011), Bailey et al. (2016), Wang and Fan (2017), Lettau and Pelger (2020), Cai et al. (2020), Bailey et al. (2021), Freyaldenhoven (2022), Uematsu and Yamagata (2022), and Pesaran and Smith (2025).
@@
In comparison to this literature, our main methodological contribution is a unified procedure for constructing confidence intervals for signal strength across all four models and all signal ranges, as presented in Section 3. This approach does not rely on standard Gaussian quantiles, but instead uses a novel random transition process $\mathcal{T}(\Theta)$, whose quantiles are tabulated in Table 1. As shown in Figure 1(a), the Gaussian approximation performs poorly near the critical threshold, making $\mathcal{T}(\Theta)$ essential for accurate inference in that regime. This is reminiscent of the construction of uniform confidence intervals for autoregressive models in Stock (1991); Mikusheva (2007), where the non-standard asymptotics near the unit root are smoothly connected to the standard normal behavior in the stationary region.
Beyond quantifying uncertainty in signal strength, our framework also enables signal detection and the assessment of factor informativeness. Specifically, one can check whether the uniform confidence intervals include zero and the identification cut-off, respectively.
@@
A surprising finding is that the same transition process $\mathcal{T}(\Theta)$ governs all four models. In fact, the proofs in Section 9 follow different paths depending on the model, and only in the final step does a structural identity emerge, revealing that all four asymptotic distributions coincide. Random matrix theory has many universality theorems, and based on our results, we predict that the same transition process $\mathcal{T}(\Theta)$ governs a much wider class of signal-plus-noise models, beyond the ones analyzed here.
@@ 1.5. Outline of the paper
Section 2 introduces the four main signal-plus-noise models. Section 3 presents a unified procedure for constructing confidence intervals for signal strengths. Section 4 lays out the theoretical foundations underlying this procedure. Section 5 offers three empirical illustrations. Some extensions are discussed in Section 6. Section 7 concludes. All proofs are in Appendices 8 and 9.
@@ 2.1. Spiked Wigner matrix
where $r$ is fixed (not growing with $N$) and $\theta_{1}>\dots>\theta_{r}>0\in\mathbb{R}$ are the strengths of $r$ signals, with corresponding directions $\mathbf{u}^{*}_{1},\dots,\mathbf{u}^{*}_{r}$, which are assumed to be orthonormal $N$–dimensional vectors. The noise matrix $\mathcal{E}$ is a (Wigner) matrix sampled from the Gaussian Orthogonal Ensemble, meaning that $\mathcal{E}=\frac{1}{\sqrt{2N}}(\mathcal{Z}+\mathcal{Z}^{\mathsf{T}})$, where $\mathcal{Z}$ is an $N\times N$ matrix of i.i.d. $\mathcal{N}(0,\sigma^{2})$ entries (see Section 6.1 for the discussion of non-Gaussian setting). We assume that $\theta_{i}$ and $\mathbf{u}^{*}_{i}$ are unknown deterministic parameters; one could alternatively allow $\mathbf{u}^{*}_{i}$ to be random, provided they are independent of $\mathcal{E}$. Our goal is to estimate the signal strengths $\theta_{1},\dots,\theta_{r}$.
We first assume that the variance of the underlying noise $\mathcal{Z}$, $\sigma^{2}$, is known and, without loss of generality, set it to $1$ by rescaling the model. [^3] The prefactor $\frac{1}{\sqrt{2N}}$ in the definition of $\mathcal{E}$ ensures that its eigenvalues remain bounded and fill the interval $[-2,2]$ as $N\to\infty$. In Section 3.3 we discuss adjustments for the case when $\sigma^{2}$ is unknown.
One common application of the spiked Wigner framework is modeling symmetric interaction networks, for instance, economic or social activity among $N$ agents. In such cases each rank‑one component $\theta_{i}\mathbf{u}^{*}_{i}(\mathbf{u}^{*}_{i})^{\top}$ captures a latent structure in agent attributes $\mathbf{u}^{*}_{i}$, while the observed interactions are contaminated by noise $\mathcal{E}$. Low-rank approximations of this form underpin seminal network models including the stochastic block model of Holland et al. (1983), where communities are inferred from block‑structured adjacency matrices, and latent space models.
@@
Informally, the proposition says that “good” recovery of $\theta_{i}$ from the largest eigenvalues of $\mathbf{A}$ is possible if and only if $\theta_{i}$ is larger than the critical value $\theta^{c}=1$. In this case, to estimate $\theta_{i}$, one should take $\lambda_{i}$ and apply the inverse of the mapping $\theta\mapsto\lambda(\theta)$, which is $\lambda\mapsto\frac{1}{2}\left(\lambda+\sqrt{\lambda^{2}-4}\right)$.
@@
(2.4) $\mathrm{Prob}\left(\theta_{i}\in[\theta^{-}_{i}(\lambda_{i},N,\alpha),\theta^{+}_{i}(\lambda_{i},N,\alpha)]\right)\approx\alpha,$
@@
We drop $o\left(\frac{1}{\sqrt{N}}\right)$ terms, plot the intervals from the preceding formula on the $(\theta,\lambda)$–plane, and then transpose the axes to obtain the desired confidence intervals on the $(\lambda,\theta)$–plane; see the 2nd method in Figure 1(a). For $\theta$ bounded away from $1$ (equivalently, $\lambda$ bounded away from $2$), this procedure is equivalent to the intervals (2.5) as $N\to\infty$, though their finite-sample behavior differs near the cutoff. Compared to Monte Carlo intervals, both methods exhibit substantial bias, but in different directions.
@@ 2.2. Spiked covariance model
For the second setup we consider a deterministic $N\times N$ matrix
@@
where $r$ is a fixed, small number, $\theta_{1}>\dots>\theta_{r}>\sigma^{2}$ are the signal strengths, and $\mathbf{u}^{*}_{1},\dots,\mathbf{u}^{*}_{r}$ are orthonormal $N$-dimensional vectors representing $r$ signal directions. The eigenvalues of $\Omega$ are $\theta_{1},\theta_{2},\dots,\theta_{r}$, and $\sigma^{2}$ with multiplicity $(N-r)$. As before, we assume that $\sigma^{2}$ is known and set it to $1$ without loss of generality; adjustments for unknown $\sigma^{2}$ are discussed in Section 3.3.
@@
As in the previous section, we can use this Gaussian approximation to construct confidence intervals for each $\theta_{i}$, yielding a modification of (2.5). However, this approach faces the same issue: the intervals become unreliable as $\lambda_{i}$ approaches $\lambda_{+}$ and must be corrected using the procedures in Section 3.
@@ 2.3. Factor model
where $r$ is a fixed small number, $\theta_{1},\dots,\theta_{r}>0$ are the signal strengths, $\mathbf{u}^{*}_{1},\dots,\mathbf{u}^{*}_{r}$ are $N$-dimensional orthonormal vectors of signal directions, called “loadings” [^6] Sometimes $\{\sqrt{\theta_{i}}\mathbf{u}^{*}_{i}\}$ rather than $\{\mathbf{u}^{*}_{i}\}$ are referred to as loadings., and $\mathbf{v}^{*}_{1},\dots,\mathbf{v}^{*}_{r}$ are $S$-dimensional orthonormal vectors called “factors”. The noise matrix $\mathcal{E}$ has independent $\mathcal{N}(0,\sigma^{2})$ entries. For now, we assume $\sigma^{2}$ to be known and set it to $1$; adjustments for unknown $\sigma^{2}$ are discussed in Section 3.3.
Our goal is to estimate $\theta_{1},\dots,\theta_{r}$ from the eigenvalues $\lambda_{1}\geq\lambda_{2}\geq\dots\geq\lambda_{N}$ of the sample covariance matrix $\frac{1}{S}XX^{\mathsf{T}}$. While the factor model has similarities to the spiked covariance model of the previous section, they are not equivalent, because we treat $\sqrt{\theta_{i}}\cdot\mathbf{u}^{*}_{i}(\mathbf{v}^{*}_{i})^{\mathsf{T}}$ in (2.10) as deterministic parameters (the models would have been equivalent up to shift $\theta_{i}\to\theta_{i}+\sigma^{2}$, if each $\sqrt{S}\mathbf{v}^{*}_{i}$ were a mean 00 Gaussian vector with i.i.d. components). This distinction allows the factor model to capture complex structures along the $S$-dimension, which is essential in applications across finance, macroeconomics, natural sciences, and other fields. Once again, an analogue of (2.3) holds.
@@
$\theta_{i}\in\left[\theta(\lambda_{i})-\frac{\sigma(\lambda_{i})}{\sqrt{N}}z_{\alpha/2},\theta(\lambda_{i})+\frac{\sigma(\lambda_{i})}{\sqrt{N}}z_{\alpha/2}\right],$
where
@@
There is also a direct analogue of the second Gaussian method described in Remark 2.1. Figure 1(b) compares these two Gaussian-based intervals with our new approach, which we present in Section 3. The comparison reveals the same key features as in the spiked Wigner model.

### v2

@@ Abstract.
The paper analyzes four classical signal plus noise models: the factor model, spiked sample covariance matrices, the sum of a Wigner matrix and a low-rank perturbation, and canonical correlation analysis with low-rank dependencies. The objective is to construct confidence intervals for the signal strength that are uniformly valid across all regimes – strong, weak, and critical signals. We demonstrate that traditional Gaussian approximations fail in the critical regime. Instead, we introduce a universal transitional distribution that enables valid inference across the entire spectrum of signal strengths. The approach is illustrated through applications in macroeconomics and finance.
The authors thank seminar participants at Cambridge, Duke, online RMTA seminar, Princeton, and Simons Symposium for their helpful comments. Gorin’s work was partially supported by NSF grant DMS-2246449. Sodin’s work was partially supported by a Philip Leverhulme Prize (PLP-2020-064) and by the Morris Belkin Visiting Professor Program at the Weizmann Institute of Science.
@@ 1.1. Motivation
In the modern era researchers increasingly have access to high-dimensional data across a wide range of fields. These data are inevitably contaminated by various forms of error and noise, making the separation of meaningful structure from background noise a central challenge. To address this analysts commonly employ dimension-reduction techniques. The two dominant approaches are low-rank methods, which assume that the underlying signal lies in a lower-dimensional subspace, and sparsity-based methods, which assume that only a small subset of variables or parameters are truly relevant, i.e., nonzero. This paper adopts the low-rank perspective. For a discussion of settings where this assumption is appropriate, we refer to Udell and Townsend (2019), Giannone et al. (2021), and Thibeault et al. (2024). In particular, Giannone et al. (2021) argue that numerous data sets in macroeconomics, microeconomics, and finance exhibit dense, rather than sparse, structures.
@@
The feasibility of consistently estimating the signal component $LF^{\mathsf{T}}$ from the observed data $X$ hinges on the strength of the signal, which can be quantified by the singular values of $LF^{\mathsf{T}}$. When these singular values are large, the signal is strong and estimation is reliable, as can be directly predicted from the form of (1.1). As the signal weakens, the data $X$ becomes less informative, and below a certain critical threshold, accurate recovery becomes impossible. The relationships between the strength of the signal and feasibility of reconstruction of $LF^{\mathsf{T}}$ have been rigorously analyzed in a number of studies, see, e.g., Stock and Watson (2002); Bai and Ng (2002); Bai (2003); Paul (2007); Onatski (2012); Johnstone and Paul (2018); Bai and Ng (2023); Fan et al. (2024b); Barigozzi and Hallin (2024) and references therein.
@@ 1.2. Models and results
We analyze four classical high-dimensional statistical models: the factor model (1.1), spiked sample covariance, the spiked Wigner model, and spiked canonical correlations. These correspond to three fundamental ensembles from random matrix theory: the Laguerre/Wishart ensemble for the first two, the Hermite/Gaussian/Wigner ensemble for the third, and the Jacobi ensemble for the fourth. Each model can be viewed as an instance of the signal plus noise framework, also known as spiked random matrices, a term originating with Johnstone (2001), in which a low-rank signal matrix is embedded in a high-dimensional noisy environment. The goal is to detect and quantify the signal.
@@
Our analysis is based on spectral methods, whereby signal strength is inferred from the eigenvalues of certain model-specific matrices. In all four setups a well-documented phase transition phenomenon arises: the signal strength can be consistently estimated (in the high-dimensional asymptotic regime with proportional growth of data dimensions) only when it exceeds a critical threshold, see Jones et al. (1978), Baik and Silverstein (2006), Onatski (2012), Bao et al. (2019) and more references in Section 2. When the signal strength falls below the threshold, only partial probabilistic information, such as asymptotics of the likelihood ratio test can be recovered, but reliable point estimation becomes impossible, see, e.g. Onatski et al. (2013, 2014); Dobriban (2017); Johnstone and Onatski (2020); El Alaoui et al. (2020). The intermediate regime, where the signal strength is close to the threshold, is typically referred to as the “critical” regime. This regime is particularly challenging for inference.
@@
Our paper fills this gap by proposing a general procedure for constructing confidence intervals for signal strength. Remarkably, across all four models we study, the confidence intervals are characterized by a common limiting (stochastic) object, which we call the Airy–Green function and denote $\mathcal{G}(w)$. Our main contributions are: a rigorous construction of this function, a unified set of theorems linking it to the four canonical models, and tabulated confidence intervals based on $\mathcal{G}(w)$. The only model-specific components are a set of scaling constants, which we provide explicitly for each setting. In addition, our results imply a formula-free, bootstrap-type procedure for constructing confidence intervals.
@@ 1.3. Econometrics and statistics contributions
In economics and finance it has long been observed that many data sets contain factors that are either non-informative or far from strong – see e.g. Giglio et al. (2023) and Kim et al. (2024) for overviews and extensive references. This concern is especially apparent in the vast “factor zoo” of potential variables proposed to explain stock returns. This empirical reality has motivated a line of theoretical research focused on inference for weaker factors. Broadly speaking, factors can be classified by their strength into three categories: strong (as in, e.g. Bai and Ng (2002); Stock and Watson (2002)), semi-strong (as in, e.g. Bai and Ng (2023); Fan et al. (2024b)), and weak [^2] What we call semi-strong factors are sometimes referred to as weak, while weak factors may be termed weakly influential or extremely weak. (as in, e.g., Onatski (2012)). The literature also includes statistical procedures for testing and distinguishing between these types of factors (see, in particular, Kim et al. (2024)). Over the past decades a growing body of research has focused specifically on factor strength, including contributions by Chudik et al. (2011), Bailey et al. (2016), Wang and Fan (2017), Lettau and Pelger (2020), Cai et al. (2020), Bailey et al. (2021), Freyaldenhoven (2022), Uematsu and Yamagata (2022), and Pesaran and Smith (2025).
@@
In comparison to this literature, our main methodological contribution is a unified procedure for constructing confidence intervals for signal strength across all four models and all signal ranges, as presented in Section 3. This approach does not rely on standard Gaussian quantiles, but instead uses a novel random transition process $\mathcal{T}(\Theta)$, whose quantiles are tabulated in Table 1. Figure 1 reveals that our procedure performs well for a wide range of different signals. The quality of approximations is excellent for critical and weak signals, and does not deteriorate for larger signals, hence, covering also the case of strong signals. In contrast, as shown in Figure 1, the Gaussian approximation performs poorly near the critical threshold, making $\mathcal{T}(\Theta)$ essential for accurate inference in that regime. Our approach is reminiscent of the construction of uniform confidence intervals for autoregressive models in Stock (1991); Mikusheva (2007), where the non-standard asymptotics near the unit root are smoothly connected to the standard normal behavior in the stationary region.
@@
A surprising finding is that the same transition process $\mathcal{T}(\Theta)$ governs all four models. In fact, the proofs in Section 9 follow different paths depending on the model, and only in the final step does a structural identity emerge, revealing that all four asymptotic distributions coincide. Random matrix theory has many universality theorems, and based on our results, we predict that the same transition process $\mathcal{T}(\Theta)$ governs a much wider class of signal plus noise models, beyond the ones analyzed here.
Beyond quantifying uncertainty in signal strength, our framework also enables signal detection and the assessment of factor informativeness. Specifically, one can check whether the uniform confidence intervals include zero and the identification threshold, respectively. Paul (2007); Onatski (2012); Benaych-Georges and Nadakuditi (2012); Bykhovskaya and Gorin (2025) show that, for weak signals, estimates of the signal direction are inconsistent. Asymptotically, the estimated direction is inclined at an angle $\phi$ relative to the true direction. There are two complementary cases: if $\phi=\frac{\pi}{2}$, then the estimated direction contains no information about the truth and can be discarded; if $\phi<\frac{\pi}{2}$, then information is present and can potentially be extracted. It turns out that $\phi$ depends on the signal strength, decreasing as the strength increases, and that the transition between these two cases occurs precisely at the identification threshold discussed above. This reinforces our results on estimating signal strength in the critical regime as a tool for distinguishing between these two cases in direction estimation.
@@ 1.5. Outline of the paper
Section 2 introduces the four main signal plus noise models. Section 3 presents a unified procedure for constructing confidence intervals for signal strengths. Section 4 lays out the theoretical foundations underlying this procedure. Section 5 offers three empirical illustrations. Extensions are discussed in Section 6. Section 7 concludes. All proofs are in Appendices 8 and 9.
@@ 2.1. Spiked Wigner matrix
where $r$ is fixed (not growing with $N$) and $\theta_{1}>\dots>\theta_{r}>0\in\mathbb{R}$ are the strengths of $r$ signals, with corresponding directions $\mathbf{u}^{*}_{1},\dots,\mathbf{u}^{*}_{r}$, which are assumed to be orthonormal NN–dimensional vectors. The noise matrix $\mathcal{E}$ is a (Wigner) matrix sampled from the Gaussian Orthogonal Ensemble, meaning that $\mathcal{E}=\frac{1}{\sqrt{2N}}(\mathcal{Z}+\mathcal{Z}^{\mathsf{T}})$, where $\mathcal{Z}$ is an $N\times N$ matrix of i.i.d. $\mathcal{N}(0,\sigma^{2})$ entries (see Section 6.1 for non-Gaussian setting). We assume that $\theta_{i}$ and $\mathbf{u}^{*}_{i}$ are unknown deterministic parameters; one could alternatively allow $\mathbf{u}^{*}_{i}$ to be random, provided they are independent of $\mathcal{E}$. Our goal is to estimate the signal strengths $\theta_{1},\dots,\theta_{r}$.
We first assume that the variance of the underlying noise $\mathcal{Z}$, $\sigma^{2}$, is known and set it to $1$ by rescaling the model. [^3] The prefactor $\frac{1}{\sqrt{2N}}$ in the definition of $\mathcal{E}$ ensures that its eigenvalues fill the interval [2,2][-2,2] as $N\to\infty$. In Section 3.4 we discuss adjustments for the case of unknown $\sigma^{2}$.
One common application of the spiked Wigner framework is modeling symmetric interaction networks, such as economic or social activity among $N$ agents. Each rank‑one component $\theta_{i}\mathbf{u}^{*}_{i}(\mathbf{u}^{*}_{i})^{\top}$ captures a latent structure in agent attributes $\mathbf{u}^{*}_{i}$, while the observed interactions are contaminated by noise $\mathcal{E}$. Low-rank approximations of this form underpin seminal network models including the stochastic block model of Holland et al. (1983), where communities are inferred from block‑structured adjacency matrices, and latent space models.
@@
Informally, the proposition says that “good” recovery of $\theta_{i}$ from the largest eigenvalues is possible if and only if $\theta_{i}$ is larger than the critical value $\theta^{c}=1$. In this case, to estimate $\theta_{i}$, one should take $\lambda_{i}$ and apply the inverse of the mapping $\theta\mapsto\lambda(\theta)$, which is $\lambda\mapsto\frac{1}{2}\left(\lambda+\sqrt{\lambda^{2}-4}\right)$.
@@
(2.4) $\mathrm{Prob}\left(\theta_{i}\in[\theta^{-}_{i}(\lambda_{i},N,\alpha),\theta^{+}_{i}(\lambda_{i},N,\alpha)]\right)\approx 1-\alpha,$
@@
We drop $o\left(\frac{1}{\sqrt{N}}\right)$ terms, plot the intervals from the preceding formula on the $(\theta,\lambda)$–plane, and then transpose the axes to obtain the desired confidence intervals on the $(\lambda,\theta)$–plane; see the 2nd method in Figure 1(a). For $\theta$ away from $1$ (equivalently, $\lambda$ bounded away from $2$), this procedure is equivalent to the intervals (2.5) as $N\to\infty$, though their finite-sample behavior differs near the cutoff. Both methods exhibit substantial bias, but in different directions.
@@ 2.2. Spiked covariance model
Second, we consider a deterministic $N\times N$ matrix
@@
where $r$ is fixed, $\theta_{1}>\dots>\theta_{r}>\sigma^{2}$ are the signal strengths, and $\mathbf{u}^{*}_{1},\dots,\mathbf{u}^{*}_{r}$ are orthonormal $N$-dimensional vectors representing $r$ signal directions. The eigenvalues of $\Omega$ are $\theta_{1},\theta_{2},\dots,\theta_{r}$, and $\sigma^{2}$ with multiplicity (Nr)(N-r). As before, we assume that $\sigma^{2}$ is known and set it to $1$ without loss of generality; adjustments for unknown $\sigma^{2}$ are discussed in Section 3.4.
@@
As in the previous section, we can use this Gaussian approximation to construct confidence intervals for each $\theta_{i}$, yielding a modification of (2.5). However, this approach faces the same issue: the intervals become unreliable as $\lambda_{i}$ approaches $\lambda_{+}$ and must be corrected.
@@ 2.3. Factor model
where $r$ is a fixed small number, $\theta_{1},\dots,\theta_{r}>0$ are the signal strengths, $\mathbf{u}^{*}_{1},\dots,\mathbf{u}^{*}_{r}$ are $N$-dimensional orthonormal vectors of signal directions, called “loadings” [^6] Sometimes $\{\sqrt{\theta_{i}}\mathbf{u}^{*}_{i}\}$ rather than $\{\mathbf{u}^{*}_{i}\}$ are referred to as loadings., and $\mathbf{v}^{*}_{1},\dots,\mathbf{v}^{*}_{r}$ are $S$-dimensional orthonormal vectors called “factors”. The noise matrix $\mathcal{E}$ has independent $\mathcal{N}(0,\sigma^{2})$ entries. For now and until Section 3.4 we assume $\sigma^{2}$ to be known and set it to $1$.
Our goal is to estimate $\theta_{1},\dots,\theta_{r}$ from the eigenvalues $\lambda_{1}\geq\lambda_{2}\geq\dots\geq\lambda_{N}$ of the sample covariance matrix $\frac{1}{S}XX^{\mathsf{T}}$. While the factor model has similarities to the spiked covariance model of the previous section, they are not equivalent, because we treat $\sqrt{\theta_{i}}\cdot\mathbf{u}^{*}_{i}(\mathbf{v}^{*}_{i})^{\mathsf{T}}$ in (2.10) as deterministic parameters (the models would have been equivalent up to shift $\theta_{i}\to\theta_{i}+\sigma^{2}$, if each $\sqrt{S}\mathbf{v}^{*}_{i}$ were a mean 0 Gaussian vector with i.i.d. components). This distinction allows the factor model to capture complex structures along the $S$-dimension, which is essential in applications across finance, macroeconomics, natural sciences, and other fields. Once again, an analogue of (2.3) holds.
@@
$\theta_{i}\in\left[\theta(\lambda_{i})-\frac{\sigma(\lambda_{i})}{\sqrt{N}}z_{\alpha/2},\theta(\lambda_{i})+\frac{\sigma(\lambda_{i})}{\sqrt{N}}z_{\alpha/2}\right],\qquad\text{ where }$
@@
There is also a direct analogue of the second Gaussian method described in Remark 2.1. Figure 1(b) compares these two Gaussian-based intervals with our new approach, presented in Section 3. The comparison reveals the same key features as in the spiked Wigner model.